# RoboRefer: Towards Spatial Referring with Reasoning in Vision-Language Models for Robotics

### v3

@@ B.1.1 Multi-Stage Image Filtering
2D Web Images aim to endow the model with basic spatial concepts and comprehensive depth perception across both indoor and outdoor scenes. Here we use OpenImagehttps://storage.googleapis.com/openimages/web/index.html [132] as 2D data source.
Overall Motivation and Goals for Filtering. The OpenImages dataset offers a vast collection of 2D internet images (1.7 M images in training split) with extensive visual diversity, but many images, such as text-only graphics, QR codes, medical scans, or abstract art, lack relevance for spatial understanding and referring with reasoning. To curate a dataset tailored for these tasks, we employ a two-stage filtering pipeline: a coarse pre-filtering using SigLIP2 [73], followed by fine-grained selection via Qwen2.5-VL [11], to retain images rich in spatial semantics.
@@ B.1.2 Pseudo-3D Scene Graphs Construction
Although filtered 2D images contain some spatial cues, QA pairs containing sufficient 3D spatial information (e.g., “near” vs. “far”, distances) derived directly from these 2D images are challenging. Inspired by prior work [1, 6], we construct pseudo-3D scene graphs from 2D images to enhance the generation of QA pairs with rich 3D spatial semantics. In these graphs, nodes represent object attributes, while edges encode inter-object spatial relations. We detail the process of converting 2D images into pseudo-3D scene graphs below.
Object Detection and Annotation. Although the OpenImages dataset provides annotations, its limited vocabulary and coarse labeling are insufficient for open-world scenarios. To address this, we leverage state-of-the-art foundation models for enhanced object detection and annotation. Specifically, our scene graph construction pipeline integrates the Recognize Anything Model (RAM) [65] and GroundingDINO [100] to assign semantic labels and bounding boxes to key objects in filtered raw 2D images. The workflow is broadly as follows:
@@
3D-aware information Extraction. To further extract 3D-aware information from 2D images, we adopt UniDepth V2 [67] for metric depth estimation due to its recent state-of-the-art performance, surpassing models such as DepthPro and Metric3Dv2 across multiple benchmarks. For camera intrinsic prediction, we employ WildeCamera [75]. Together, these models enable robust 3D point cloud reconstruction of the scene. Based on previously annotated object bounding boxes, we apply SAM 2.1 [77] to generate instance masks. Each resulting Pseudo-3D scene graph comprises object labels (via RAM), 2D bounding boxes (via GroundingDINO), instance masks (via SAM 2.1), and object-level point clouds (via UniDepth v2 with WildeCamera), resulting in axis-aligned 3D bounding boxes. A visualization is provided in Fig. 8.
@@ B.1.3 Hierarchical Object Description Generation
While 3D scene graphs encode basic object categories, real-world scenes often contain multiple instances per category. Prior datasets [3, 2] simplify reference by assuming a single object per category, limiting their utility in spatial referring tasks. To overcome this, we augment object descriptions with attributes and spatial relations, enabling finer-grained disambiguation among similar instances of the same category. We present our two-stage generation pipeline below.
@@ B.1.4 Generating Diverse QA Pairs via Pseudo-3D Scene Graphs
Reasoning QA Generation. To generate more natural, complex, and diverse QA pairs beyond templated formats, we leverage a powerful reasoning LLM, QwQ-32B [143]. It takes the factual statements, initial QA pairs, and multiple-choice questions (if available) as input, along with global image captions and precise object descriptions. QwQ-32B then produces more challenging and conversational spatial reasoning QA. The prompt design is shown in Listing 8.
@@ B.2.1 Why Use CA-1M and How to Pre-process It
Rationale for Selecting CA-1M as the 3D Data Source. To enable fine-grained spatial reasoning in indoor environments, we adopt Apple’s open-source CA-1Mhttps://github.com/apple/ml-cubifyanything [133] dataset as our primary 3D data source. CA-1M aligns closely with our objectives due to the following key attributes:
@@
    ARKitScenes [144]: As a predecessor to CA-1M, it provides only global 3D bounding boxes without per-frame 3D or 2D annotations. Projecting these 3D boxes to 2D yields oversized and contain irrelevant objects. Additionally, its annotations are less comprehensive, and image resolution is also much lower compared to CA-1M.
@@
    ScanNet V2 [145]: Lacks 3D bounding boxes, making object orientation estimation infeasible. Although EmbodiedScan introduces per-frame 3D bounding boxes, it still lacks 2D annotations, and projected 2D boxes remain imprecise and contain irrelevant objects.
@@
    3RScan [146]: Suffers from low image quality, hindering spatial information extraction and limiting its usability.
@@ B.3.1 Indoor Scene Generation
Initial Scene Generation. We utilize Infinigen [134] to generate a large corpus of indoor scenes. To be specific, we configure the generation process to exclude small objects by setting compose_indoors.solve_steps_small=0 to avoid pre-existing clutter on target surfaces. This allows us to reserve space for the subsequent placement of our curated 3D assets. This initial step yields over 3k unique indoor scenes.
@@
Our 3D assets are sourced from the Objaverse [135] LVIS dataset and undergo a two-stage filtering process to ensure quality and relevance.
@@
Stage 2: Attribute-based Filtering. Next, we apply fine-grained filtering using attributes from the OrienText300K [7] dataset. Retained assets satisfy the following criteria:
@@ D.1 Architecture
We adopt NVILA [37] as base model, including a visual encoder, an LLM, and a multimodal projector.
Visual Encoder. We use the same image encoder as siglip-so400m-patch14-448 [74] from NVILA [37], supporting $448\times 448$ resolution for richer visual details. Rather than simply resizing the image to a fixed resolution and producing the same number of tokens, this image encoder processes inputs at dynamic resolutions, yielding more visual tokens from higher-resolution images via finer patch division. This enables fine-grained vision-language understanding, crucial for tasks like point prediction that require detailed perception beyond VQA. We further incorporate a dedicated depth encoder, structurally mirroring the image encoder and initialized with its weights. It encodes relative depth maps as special images, providing spatial cues to enhance 3D understanding.
Large Language Model. We adopt the Qwen2 LLM backbone from NVILA [37], which has been fully fine-tuned with extensive data during supervised training. This endows the model with rich visual knowledge, facilitating downstream 3D spatial understanding and reasoning tasks.
Multi-Modal Projector. To align multi-modal representations (e.g., image to language, depth to language), we use linear connectors, the same as NVILA [37], which is better than Q-Former, to allow the LLM to focus on visual understanding and improve generalization. Separate connectors for image and depth embeddings ensure modality-specific processing and prevent cross-modal interference.
@@ D.2 Training Data
SFT stage. Specifically, in the first step of the SFT stage, i.e., depth alignment, we train a depth projector to align depth and language space using the RefSpatial (RGB-D) dataset with 2.52.5M samples. To increase training efficiency, we slice multi-turn conversations (up to $15$ turns per sample), yielding 3.43.4M samples post-processing to train our model. In the second step, i.e., spatial understanding enhancement via full-parameter fine-tuning—we use both RefSpatial (RGB) and RefSpatial (RGB-D) datasets, yielding 6.86.8M samples after slicing. To further improve instruction-following and referring capabilities, we incorporate auxiliary datasets: 965965k samples from instruction-tuned data (LLaVA-1.5 [131], LRV [130]), 321321k from referring datasets (RefCOCO/+/g [86]), 176176k from SAT [4] benchmark training sets, and 127127k from EmbSpatial [22] benchmark training sets. These additions help bridge distribution gaps between RefSpatial and benchmark-style queries. After slicing, the total number of samples used in this stage reaches 8.58.5M post-slicing.
@@
SFT (S.U.E) Spatial RefSpatial (RGB), RefSpatial (RGB-D), SAT [4], EmbSpatial [22]
General COCO [147], GQA [18], OCR-VQA [148], TextVQA [149], VG [150], LRV [130]
REC RefCOCO/+/g [86]
@@ D.3 SFT Training Details
To be specific, our SFT consists of two steps. In the first step, depth alignment, only the depth projector is updated by using the RefSpatial (RGB-D). We employ a maximum learning rate of 1e-4, a weight decay of 0, and a warm-up ratio of 0.03. The 2B variant is trained with a batch size of 7 per GPU, and the 8B variant with 3, both for one epoch. In the second step of spatial understanding enhancement, we fine-tune all model parameters using the datasets described in Sec. D.2. Training is conducted for one epoch with a maximum learning rate of 5e-5. We use a batch size of 6 per GPU for the 2B model and 2 for the 8B model. Other hyperparameters follow those in the first step. For more details, please refer to NVILA [37] settings during alignment and SFT.
@@ D.4 RFT Training Details
During the RFT stage, we refine $\pi_{\mathrm{SFT}}$ via GRPO [113], a reinforcement learning method designed for efficiency and scalability. Unlike PPO [151], which relies on a costly value network, GRPO estimates relative advantages by comparing intra-group rewards, reducing computation, and simplifying optimization. This makes it well-suited for reasoning-intensive spatial referring tasks. In detail, we modify R1-V [121] to support our 3D-aware architecture. Training is conducted for two epochs with a batch size of 1 per GPU and 8 outputs in GRPO. For details about hyperparameters, see R1-V [121].
@@ D.4.2 Reward Design and Policy Update
Point L1 Reward $R_{P}$. This component evaluates the accuracy of the model’s final point prediction by comparing it with the ground truth from the annotations of RefSpatial. Following the criterion inspired by Seg-zero [128], a stricter reward of 1 is assigned if the L1 distance between the predicted and ground-truth points is within 50 pixels; otherwise, the reward is 0.
Notably, most process-based rewards depend on a Process Reward Model (PRM), typically a fine-tuned LLM or VLM tasked with providing feedback. However, applying such an approach in our setting presents two main challenges. (1) LLMs cannot process images, making it impossible to determine whether predicted coordinates match the target object. (2) Although VLMs integrate visual and textual information, prior work [152] has shown they may lack precise visual understanding when dealing with textual coordinates. Since the correct assessment of predicted coordinates is paramount for reward assignment, additional or specialized methods are needed to ensure reliable feedback.
To address this issue, we propose a rule-based process reward for spatial referring that obviates the need for a Process Reward Model. Our approach directly evaluates key intermediate perceptual steps using the ground-truth step-wise annotations provided in RefSpatial. This contrasts with most existing methods on process-based rewards [153, 154], which emphasize strictly sequential reasoning and rely on a PRM for evaluation. In contrast, our method employs metric-sensitive rule-based process reward functions to assess intermediate perceptual results in an order-invariant manner. Our key insight lies in two aspects: (1) Metric-sensitivity: Different spatial attributes require distinct metrics due to inherent differences in their representations (e.g., points for positions, vectors for orientations). (2) Order-invariance: The reasoning process in spatial referring is not strictly sequential; for instance, identifying the position of the keyboard or the mouse first does not affect the final interpretation of “the free area between the keyboard and the mouse”.
@@ E.1 Experiments Compute Resources
Spatial Referring in RFT. The process is conducted on 1 node over 3 days for 2B variants. However, our model is over twice as slow as other Qwen 2/2.5-VL-based methods [121, 127], mainly because they process only a single RGB image during training and can leverage vLLM for group inference acceleration. In contrast, our method requires RGB-D inputs and modifies the original NVILA architecture, making it incompatible with vLLM or SGLang acceleration.
@@ E.2 Spatial Understanding Benchmarks
We evaluate several public single-step spatial understanding benchmarks, including CV-Bench [15] (2D Spatial Relation, 3D Depth Order, 3D Distance), the BLINK [16] validation set (Spatial Relation, Relative Depth), RoboSpatial [2] (configuration), SAT [4], and EmbSpatial [22], following their official evaluation protocols. We exclude non-spatial tasks from our evaluation, such as 2D Counting in CV-Bench and Art Style or IQ Test in BLINK. Since all these benchmarks are multiple-choice tasks, we report accuracy as the evaluation metric.
We compare $3$ categories of models: (1) proprietary VLMs, such as Gemini-2.5-Pro [9], which show strong spatial perception, as shown in the Gemini-Robotics [106] paper; (2) open-source VLMs trained on general VQA datasets; and (3) spatially specialized models trained on spatially relevant datasets, offering basic spatial understanding.

### v4

@@ B.1.1 Multi-Stage Image Filtering
2D Web Images aim to endow the model with basic spatial concepts and comprehensive depth perception across both indoor and outdoor scenes. Here we use OpenImagehttps://storage.googleapis.com/openimages/web/index.html [kuznetsova2020open] as 2D data source.
Overall Motivation and Goals for Filtering. The OpenImages dataset offers a vast collection of 2D internet images (1.7 M images in training split) with extensive visual diversity, but many images, such as text-only graphics, QR codes, medical scans, or abstract art, lack relevance for spatial understanding and referring with reasoning. To curate a dataset tailored for these tasks, we employ a two-stage filtering pipeline: a coarse pre-filtering using SigLIP2 [tschannen2025siglip], followed by fine-grained selection via Qwen2.5-VL [bai2025qwen2], to retain images rich in spatial semantics.
@@ B.1.2 Pseudo-3D Scene Graphs Construction
Although filtered 2D images contain some spatial cues, QA pairs containing sufficient 3D spatial information (e.g., “near” vs. “far”, distances) derived directly from these 2D images are challenging. Inspired by prior work [cheng2024spatialrgpt, chen2024spatialvlm], we construct pseudo-3D scene graphs from 2D images to enhance the generation of QA pairs with rich 3D spatial semantics. In these graphs, nodes represent object attributes, while edges encode inter-object spatial relations. We detail the process of converting 2D images into pseudo-3D scene graphs below.
Object Detection and Annotation. Although the OpenImages dataset provides annotations, its limited vocabulary and coarse labeling are insufficient for open-world scenarios. To address this, we leverage state-of-the-art foundation models for enhanced object detection and annotation. Specifically, our scene graph construction pipeline integrates the Recognize Anything Model (RAM) [zhang2024recognize] and GroundingDINO [liu2024grounding] to assign semantic labels and bounding boxes to key objects in filtered raw 2D images. The workflow is broadly as follows:
@@
3D-aware information Extraction. To further extract 3D-aware information from 2D images, we adopt UniDepth V2 [piccinelli2025unidepthv2] for metric depth estimation due to its recent state-of-the-art performance, surpassing models such as DepthPro and Metric3Dv2 across multiple benchmarks. For camera intrinsic prediction, we employ WildeCamera [zhu2023tame]. Together, these models enable robust 3D point cloud reconstruction of the scene. Based on previously annotated object bounding boxes, we apply SAM 2.1 [ravi2024sam] to generate instance masks. Each resulting Pseudo-3D scene graph comprises object labels (via RAM), 2D bounding boxes (via GroundingDINO), instance masks (via SAM 2.1), and object-level point clouds (via UniDepth v2 with WildeCamera), resulting in axis-aligned 3D bounding boxes. A visualization is provided in Fig. 8.
@@ B.1.3 Hierarchical Object Description Generation
While 3D scene graphs encode basic object categories, real-world scenes often contain multiple instances per category. Prior datasets [daxberger2025mm, song2024robospatial] simplify reference by assuming a single object per category, limiting their utility in spatial referring tasks. To overcome this, we augment object descriptions with attributes and spatial relations, enabling finer-grained disambiguation among similar instances of the same category. We present our two-stage generation pipeline below.
@@ B.1.4 Generating Diverse QA Pairs via Pseudo-3D Scene Graphs
Reasoning QA Generation. To generate more natural, complex, and diverse QA pairs beyond templated formats, we leverage a powerful reasoning LLM, QwQ-32B [qwq32b]. It takes the factual statements, initial QA pairs, and multiple-choice questions (if available) as input, along with global image captions and precise object descriptions. QwQ-32B then produces more challenging and conversational spatial reasoning QA. The prompt design is shown in Listing 8.
@@ B.2.1 Why Use CA-1M and How to Pre-process It
Rationale for Selecting CA-1M as the 3D Data Source. To enable fine-grained spatial reasoning in indoor environments, we adopt Apple’s open-source CA-1Mhttps://github.com/apple/ml-cubifyanything [lazarow2024cubify] dataset as our primary 3D data source. CA-1M aligns closely with our objectives due to the following key attributes:
@@
    ARKitScenes [baruch2021arkitscenes]: As a predecessor to CA-1M, it provides only global 3D bounding boxes without per-frame 3D or 2D annotations. Projecting these 3D boxes to 2D yields oversized and contain irrelevant objects. Additionally, its annotations are less comprehensive, and image resolution is also much lower compared to CA-1M.
@@
    ScanNet V2 [dai2017scannet]: Lacks 3D bounding boxes, making object orientation estimation infeasible. Although EmbodiedScan introduces per-frame 3D bounding boxes, it still lacks 2D annotations, and projected 2D boxes remain imprecise and contain irrelevant objects.
@@
    3RScan [wald2020beyond]: Suffers from low image quality, hindering spatial information extraction and limiting its usability.
@@ B.3.1 Indoor Scene Generation
Initial Scene Generation. We utilize Infinigen [raistrick2024infinigen] to generate a large corpus of indoor scenes. To be specific, we configure the generation process to exclude small objects by setting compose_indoors.solve_steps_small=0 to avoid pre-existing clutter on target surfaces. This allows us to reserve space for the subsequent placement of our curated 3D assets. This initial step yields over 3k unique indoor scenes.
@@
Our 3D assets are sourced from the Objaverse [deitke2023objaverse] LVIS dataset and undergo a two-stage filtering process to ensure quality and relevance.
@@
Stage 2: Attribute-based Filtering. Next, we apply fine-grained filtering using attributes from the OrienText300K [qi2025sofar] dataset. Retained assets satisfy the following criteria:
@@ D.1 Architecture
We adopt NVILA [liu2024nvila] as base model, including a visual encoder, an LLM, and a multimodal projector.
Visual Encoder. We use the same image encoder as siglip-so400m-patch14-448 [zhai2023sigmoid] from NVILA [liu2024nvila], supporting $448\times 448$ resolution for richer visual details. Rather than simply resizing the image to a fixed resolution and producing the same number of tokens, this image encoder processes inputs at dynamic resolutions, yielding more visual tokens from higher-resolution images via finer patch division. This enables fine-grained vision-language understanding, crucial for tasks like point prediction that require detailed perception beyond VQA. We further incorporate a dedicated depth encoder, structurally mirroring the image encoder and initialized with its weights. It encodes relative depth maps as special images, providing spatial cues to enhance 3D understanding.
Large Language Model. We adopt the Qwen2 LLM backbone from NVILA [liu2024nvila], which has been fully fine-tuned with extensive data during supervised training. This endows the model with rich visual knowledge, facilitating downstream 3D spatial understanding and reasoning tasks.
Multi-Modal Projector. To align multi-modal representations (e.g., image to language, depth to language), we use linear connectors, the same as NVILA [liu2024nvila], which is better than Q-Former, to allow the LLM to focus on visual understanding and improve generalization. Separate connectors for image and depth embeddings ensure modality-specific processing and prevent cross-modal interference.
@@ D.2 Training Data
SFT stage. Specifically, in the first step of the SFT stage, i.e., depth alignment, we train a depth projector to align depth and language space using the RefSpatial (RGB-D) dataset with 2.52.5M samples. To increase training efficiency, we slice multi-turn conversations (up to $15$ turns per sample), yielding 3.43.4M samples post-processing to train our model. In the second step, i.e., spatial understanding enhancement via full-parameter fine-tuning—we use both RefSpatial (RGB) and RefSpatial (RGB-D) datasets, yielding 6.86.8M samples after slicing. To further improve instruction-following and referring capabilities, we incorporate auxiliary datasets: 965965k samples from instruction-tuned data (LLaVA-1.5 [liu2024improved], LRV [liu2024mitigating]), 321321k from referring datasets (RefCOCO/+/g [yu2016modeling]), 176176k from SAT [ray2024sat] benchmark training sets, and 127127k from EmbSpatial [du2024embspatial] benchmark training sets. These additions help bridge distribution gaps between RefSpatial and benchmark-style queries. After slicing, the total number of samples used in this stage reaches 8.58.5M post-slicing.
@@
SFT (S.U.E) Spatial RefSpatial (RGB), RefSpatial (RGB-D), SAT [ray2024sat], EmbSpatial [du2024embspatial]
General COCO [lin2014microsoft], GQA [hudson2019gqa], OCR-VQA [mishra2019ocr], TextVQA [singh2019towards], VG [krishna2017visual], LRV [liu2024mitigating]
REC RefCOCO/+/g [yu2016modeling]
@@ D.3 SFT Training Details
To be specific, our SFT consists of two steps. In the first step, depth alignment, only the depth projector is updated by using the RefSpatial (RGB-D). We employ a maximum learning rate of 1e-4, a weight decay of 0, and a warm-up ratio of 0.03. The 2B variant is trained with a batch size of 7 per GPU, and the 8B variant with 3, both for one epoch. In the second step of spatial understanding enhancement, we fine-tune all model parameters using the datasets described in Sec. D.2. Training is conducted for one epoch with a maximum learning rate of 5e-5. We use a batch size of 6 per GPU for the 2B model and 2 for the 8B model. Other hyperparameters follow those in the first step. For more details, please refer to NVILA [liu2024nvila] settings during alignment and SFT.
@@ D.4 RFT Training Details
During the RFT stage, we refine $\pi_{\mathrm{SFT}}$ via GRPO [shao2024deepseekmath], a reinforcement learning method designed for efficiency and scalability. Unlike PPO [schulman2017proximal], which relies on a costly value network, GRPO estimates relative advantages by comparing intra-group rewards, reducing computation, and simplifying optimization. This makes it well-suited for reasoning-intensive spatial referring tasks. In detail, we modify R1-V [zhang2025r1] to support our 3D-aware architecture. Training is conducted for two epochs with a batch size of 1 per GPU and 8 outputs in GRPO. For details about hyperparameters, see R1-V [zhang2025r1].
@@ D.4.2 Reward Design and Policy Update
Point L1 Reward $R_{P}$. This component evaluates the accuracy of the model’s final point prediction by comparing it with the ground truth from the annotations of RefSpatial. Following the criterion inspired by Seg-zero [liu2025seg], a stricter reward of 1 is assigned if the L1 distance between the predicted and ground-truth points is within 50 pixels; otherwise, the reward is 0.
Notably, most process-based rewards depend on a Process Reward Model (PRM), typically a fine-tuned LLM or VLM tasked with providing feedback. However, applying such an approach in our setting presents two main challenges. (1) LLMs cannot process images, making it impossible to determine whether predicted coordinates match the target object. (2) Although VLMs integrate visual and textual information, prior work [majumdar2024openeqa] has shown they may lack precise visual understanding when dealing with textual coordinates. Since the correct assessment of predicted coordinates is paramount for reward assignment, additional or specialized methods are needed to ensure reliable feedback.
To address this issue, we propose a rule-based process reward for spatial referring that obviates the need for a Process Reward Model. Our approach directly evaluates key intermediate perceptual steps using the ground-truth step-wise annotations provided in RefSpatial. This contrasts with most existing methods on process-based rewards [khalifa2025process, liu2025can], which emphasize strictly sequential reasoning and rely on a PRM for evaluation. In contrast, our method employs metric-sensitive rule-based process reward functions to assess intermediate perceptual results in an order-invariant manner. Our key insight lies in two aspects: (1) Metric-sensitivity: Different spatial attributes require distinct metrics due to inherent differences in their representations (e.g., points for positions, vectors for orientations). (2) Order-invariance: The reasoning process in spatial referring is not strictly sequential; for instance, identifying the position of the keyboard or the mouse first does not affect the final interpretation of “the free area between the keyboard and the mouse”.
@@ E.1 Experiments Compute Resources
Spatial Referring in RFT. The process is conducted on 1 node over 3 days for 2B variants. However, our model is over twice as slow as other Qwen 2/2.5-VL-based methods [zhang2025r1, shen2025vlm], mainly because they process only a single RGB image during training and can leverage vLLM for group inference acceleration. In contrast, our method requires RGB-D inputs and modifies the original NVILA architecture, making it incompatible with vLLM or SGLang acceleration.
@@ E.2 Spatial Understanding Benchmarks
We evaluate several public single-step spatial understanding benchmarks, including CV-Bench [tong2024cambrian] (2D Spatial Relation, 3D Depth Order, 3D Distance), the BLINK [fu2024blink] validation set (Spatial Relation, Relative Depth), RoboSpatial [song2024robospatial] (configuration), SAT [ray2024sat], and EmbSpatial [du2024embspatial], following their official evaluation protocols. We exclude non-spatial tasks from our evaluation, such as 2D Counting in CV-Bench and Art Style or IQ Test in BLINK. Since all these benchmarks are multiple-choice tasks, we report accuracy as the evaluation metric.
We compare $3$ categories of models: (1) proprietary VLMs, such as Gemini-2.5-Pro [team2023gemini], which show strong spatial perception, as shown in the Gemini-Robotics [team2025gemini] paper; (2) open-source VLMs trained on general VQA datasets; and (3) spatially specialized models trained on spatially relevant datasets, offering basic spatial understanding.